# Joint-Embedding Masked Autoencoder for Self-supervised Learning of Dynamic Functional Connectivity from the Human Brain

### v1

@@ Abstract
Graph Neural Networks (GNNs) have shown promise in learning dynamic functional connectivity for distinguishing phenotypes from human brain networks. However, obtaining extensive labeled clinical data for training is often resource-intensive, making practical application difficult. Leveraging unlabeled data thus becomes crucial for representation learning in a label-scarce setting. Although generative self-supervised learning techniques, especially masked autoencoders, have shown promising results in representation learning in various domains, their application to dynamic graphs for dynamic functional connectivity remains underexplored, facing challenges in capturing high-level semantic representations. Here, we introduce the Spatio-Temporal Joint Embedding Masked Autoencoder (ST-JEMA), drawing inspiration from the Joint Embedding Predictive Architecture (JEPA) in computer vision. ST-JEMA employs a JEPA-inspired strategy for reconstructing dynamic graphs, which enables the learning of higher-level semantic representations considering temporal perspectives, addressing the challenges in fMRI data representation learning. Utilizing the large-scale UK Biobank dataset for self-supervised learning, ST-JEMA shows exceptional representation learning performance on dynamic functional connectivity demonstrating superiority over previous methods in predicting phenotypes and psychiatric diagnoses across eight benchmark fMRI datasets even with limited samples and effectiveness of temporal reconstruction on missing data scenarios. These findings highlight the potential of our approach as a robust representation learning method for leveraging label-scarce fMRI data.
Index Terms:
Functional connectivity, Dynamic graph, Spatio-temporal graph, Self-supervised learning, Representation learning
I Introduction
Recent advancements in neuroimaging research have shown a growing interest in leveraging functional Magnetic Resonance Imaging (fMRI), which captures the Blood-Oxygen-Level-Dependent (BOLD) signal to understand the intricate functionalities of the human brain [16, 37]. Consequently, leveraging fMRI data through Functional Connectivity (FC) has gained traction in addressing various human brain-related problems using neural network models [18]. In this context, FC refers to the temporal correlation degree between Regions of Interest (ROI) s of the brain, enabling the mathematical interpretation of fMRI data as graph structures, where nodes represent ROI s and edges represent the connectivity or correlation between these regions. This enables the utilization of graph-based learning methods such as Graph Neural Network (GNN) to analyze and interpret brain dynamics [11, 17].
@@ I Introduction
In response, utilizing unlabeled fMRI data through Self-Supervised Learning (SSL) presents a promising solution to extract valuable features for downstream tasks [38, 34, 14]. SSL, focused on pre-training backbone models without labels, has driven performance enhancements in diverse domains, including Natural Language Processing (NLP) [10], Image [6, 13], and Video Processing [32]. Among various concepts of SSL methods, generative SSL methods inspired by Masked Language Modeling (MLM) approach [10] have demonstrated superior performance compared to the contrastive learning methods in downstream tasks [13, 32] across a range of different tasks.
@@
This trend towards generative approaches in SSL is also reflected in graph data. GraphMAE [14], for instance, applies a similar philosophy to static graphs, demonstrating that generative SSL can effectively learn foundational models in static graph contexts, surpassing the performance of traditional contrastive learning-based SSL methods. In the context of dynamic graphs, Choi et al. [7] have developed a generative SSL framework that enables foundational models to effectively capture temporal dynamics. When applied to fMRI datasets, this framework demonstrates enhanced performance across multiple downstream tasks, surpassing models based on static graph methodologies and contrastive learning approaches.
@@
While recent advancements in generative SSL have shown improved performance in learning graph data structures, existing methods tend to focus on reconstructing lower-level features [14, 7], potentially limiting their ability to grasp higher-level semantic representations. This learning of low-level features prevents the foundation model from learning the general information required for various downstream tasks in the upstream dataset, which hinders the overall performance after fine-tuning.
@@
To tackle this limitation of current generative SSL methods — reconstructing low-level representation in temporal graph data — we propose a novel framework called Spatio-Temporal Joint Embedding Masked Autoencoder (ST-JEMA) for dynamic graphs such as dynamic functional connectivity of fMRI data. In ST-JEMA, we propose a new joint embedding reconstruction loss based on the latest SSL frameworks known as Joint-Embedding Predictive Architecture (JEPA) [2]. Based on the JEPA-inspired loss function, our method induces the encoder to focus on reconstructing the target input representations rather than the target input features, across both spatial and temporal dimensions. This encourages the GNN models to learn more advanced representations of dynamic graphs.
@@
We employed the extensive UK Biobank fMRI dataset, which consists of 40,913 records, for comparing various SSL methods without demographic characteristics (e.g. Gender), and validated our methodology by fine-tuning across eight benchmarks including non-clinical and clinical fMRI datasets. The experimental results demonstrated that ST-JEMA, by considering temporal dynamics, effectively enhances the learning of node representations in the GNN encoder compared to other baseline SSL methods. This improvement was evident in its superior performance on downstream fMRI tasks.
Figure 1: Overview of Spatio-Temporal Joint Embedding Masked Autoencoder framework.
@@
  • We introduce ST-JEMA, a novel approach that overcomes the focus on low-level representations in generative SSL by enabling higher-level representation learning in dynamic graphs through JEPA-based reconstruction.
@@
II Related Works
II-A Learning Functional Connectivity of the Brain with Graph Neural Networks
Recent advancements in GNNs have significant interest in understanding brain networks, employing FC networks from fMRI data. Pioneering works have demonstrated the potential of GNNs in enhancing our understanding of functional brain dynamics by addressing the challenges of analyzing large dimensional activation of brain data [36], leveraging ROI-aware GNN architectures [21], and providing generative frameworks [15].
@@ II-A Learning Functional Connectivity of the Brain with Graph Neural Networks
II-B Self-supervised Learning on Graph Data
Utilizing SSL on graph data has proven to be an effective approach for extracting meaningful representations from large amounts of unlabeled graph data. In the realm of graph contrastive SSL, Deep Graph Infomax [30] and GraphCL [38] have been pioneering efforts that leverage augmentations and multiple graph views to enrich representation learning. However, challenges in augmentation efficacy across domains have led to innovative works, which simplify the process by introducing perturbations to embeddings [39] or model parameters [34]. Generative approaches in graph self-supervised learning focus on reconstructing the original graph structure or node features from encoded representations to capture inherent structural relationships. VGAE [19] employs variational autoencoders for adjacency matrix reconstruction enhancing link prediction performance, while GraphMAE [14] employs a masked autoencoder approach, which shows enhanced performance in node and graph classification compared to previous contrastive methods.
@@ II-B Self-supervised Learning on Graph Data
II-C Joint Embedding Predictive Architecture
@@
In the evolving landscape of SSL in computer vision, Assran et al. [2] introduces the JEPA, a new framework that integrates the concepts of both Joint Embedding Architecture (JEA) in contrastive learning and generative learning process of the Masked Autoencoder [MAE; 13]. This approach diverges from the MAE, which mainly focuses on pixel-level image reconstruction. Unlike MAE, which optimizes a decoder to reconstruct parts of masked target image patches from unmasked context patch representations, JEPA aims to avoid the potential pitfall of generating representations overly fixated on low-level details. Instead, Assran et al. [2] utilizes a predictive model trained to infer the representations of masked target patches based on unmasked context patches, enabling the encoder to formulate representations that encapsulate higher-level semantic information beyond mere pixel details. Refer to Appendix Appendix F for more detailed discussions on the related works.
III Preliminaries
III-A Problem Definition
Our goal is to develop a SSL framework capable of effectively capturing the temporal dynamics inherent in dynamic graphs derived from large-scale, unlabeled fMRI data. To achieve this, we utilize SSL strategies to train a GNN encoder and decoder, respectively denoted as $f(\cdot;{\boldsymbol{\theta}})$ and $g(\cdot;{\boldsymbol{\phi}})$, where ${\boldsymbol{\theta}}$ and ${\boldsymbol{\phi}}$ represent the model parameters. The objective is to obtain a pre-trained encoder, $f(\cdot;{\boldsymbol{\theta}}^{*})$, with optimized parameters ${\boldsymbol{\theta}}^{*}$ through SSL using the upstream dataset. Note that the decoder will only be used in pre-training and will be discarded for the finetuning. Still, carefully selecting an appropriate architecture for the decoder remains crucial for effective pre-training of the encoder.
After the pre-training phase, we fine-tune the pre-trained GNN encoder, adding specific prediction heads $h(\cdot;{\boldsymbol{\omega}})$ to solve the tasks on downstream datasets. Given the limited number of labeled data for the downstream training dataset, it is crucial to ensure that the encoder and prediction head, avoid overfit to the downstream data and achieve good generalization performance to unseen datasets. The key lies in finding a well-pre-trained encoder $f(\cdot;{\boldsymbol{\theta}}^{*})$ that captures the key patterns and dynamics within the fMRI data, which facilitates a comprehensive understanding of the fMRI dataset.
III-B Notations for Dynamic Graphs
A graph ${\mathcal{G}}$ is defined as ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$, where ${\mathcal{V}}$ denotes a set of vertices or nodes, and ${\mathcal{E}}$ denotes a set of edges or links between these vertices. The set ${\mathcal{E}}$ is defined as =${\mathcal{E}}=\{e_{ij}\,|\,i,j\in\{1,\cdots,N\},\text{ and }j\in{\mathcal{N}}(i)\}$, where $N$ is the number of nodes $|{\mathcal{V}}|$ and ${\mathcal{N}}(i)$ is the neighborhood of node $i$. Here, $e_{ij}$ indicates a relationship or connection between node $i$ and node $j$. Each node in ${\mathcal{V}}$ represents an individual entity, characterized by its feature vector $\boldsymbol{x}\in\mathbb{R}^{d_{v}}$, where $d_{v}$ indicates the dimensionality of the node feature space.
Additionally, the adjacency matrix $\boldsymbol{A}$ plays a crucial role as an edge feature in GNN. It represents the presence or absence of connections between nodes, thereby facilitating message passing that considers the relationships between nodes. Mathematically, the adjacency matrix is defined as
$\boldsymbol{A}=(a_{i,j})\in\{0,1\}^{N\times N}.$ (1)
Here, $a_{ij}=1$ when $e_{i,j}\in{\mathcal{E}}$, and otherwise 0 for all $i,j\in[N]$.
Dynamic Graph. In real-world graph datasets, there are two types of datasets: 1) static graph datasets and 2) dynamic graph datasets. Static graph datasets contain time-independent graph data that remains unchanged over time. In contrast, dynamic graph datasets include time-dependent graph data that evolves as time progresses. For dynamic graphs, it becomes crucial to consider the dynamic nature when examining numerous real-world graph data scenarios, particularly in the context of studying the functionalities of the human brain [17]. Here, a dynamic graph can be denoted as a series of evolving graphs ${\mathcal{G}}(t)=({\mathcal{V}}(t),{\mathcal{E}}(t))$ over discrete time points $t\in{\mathcal{T}}$. Each ${\mathcal{G}}(t)$ also include node features $\boldsymbol{X}(t)\in\mathbb{R}^{N\times d_{v}}$ and an adjacency matrix $\boldsymbol{A}(t)$ corresponding to the time point $t$.
III-C Constructing Dynamic Graphs from fMRI data
@@
III-D Context and Target for Reconstruction
Our method expands generative SSL upon the idea of JEPA, which reconstructs the latent representation space instead of the input feature space, within the domain of dynamic graphs derived from fMRI data. Assran et al. [2] employs a multi-block masking strategy for target block prediction, where it simultaneously utilizes $K$ different masks for a single image data. Adapting JEPA concepts to dynamic graph scenarios involves treating each masked section of the graph as targets while considering the remaining areas as contexts. Within a graph ${\mathcal{G}}(t)$ at a specific time $t$, we distinguish between masked target nodes and unmasked context nodes as follows:
@@ III-D Context and Target for Reconstruction
where $k\in[K]$ and $\boldsymbol{M}_{\boldsymbol{X}}^{(k)}(t)=\{0,1\}^{\mathbb{R}^{N\times d_{v}}}$ is the $k$-th binary mask with the same shape as $\boldsymbol{X}(t)$ at time $t\in[T_{{\mathcal{G}}}]$. Here, $\mathbf{1}_{d_{v}}$ represents a vector of ones, and $\mathbf{0}_{d_{v}}$ represents a vector of zeros, indicating masked and unmasked node features, respectively.
Similarly, the adjacency matrix $\boldsymbol{A}(t)$ is partitioned into context and target components. The context parts of the adjacency matrix are used to extract node embeddings of GNN in combination with the context node features. These embeddings are then applied to reconstruct each target component of the adjacency matrix.
IV Methods
IV-A Joint Embedding Architecture
@@
Employing separate encoders for context and target components is crucial to prevent representation collapse [12] during training [2]. Although these encoders share identical structural designs, they follow distinct parameter update rules. Specifically, the context encoder $f_{\texttt{cxt}}$ is updated via Stochastic Gradient Descent (SGD), driven by subsequent reconstruction loss objectives, while the target encoder $f_{\texttt{tar}}$ remains frozen. Subsequently, The parameters of target encoder, $\bar{{\boldsymbol{\theta}}}$, are then updated using Exponential Moving Average (EMA) based on the evolving parameters of $f_{\texttt{cxt}}$ as follows:
@@
where hyperparameter $\beta\in[0,1]$ represents the decay factor. Here, $\bar{{\boldsymbol{\theta}}}_{\text{old}}$ and $\bar{{\boldsymbol{\theta}}}_{\text{new}}$ refer to the parameters of $f_{\texttt{tar}}$ before and after applying EMA process, respectively, and ${\boldsymbol{\theta}}_{\text{ctx}}$ indicates the current parameters of $f_{\texttt{cxt}}$. Adopting distinct update rules for each encoder, as outlined, serves to prevent the encoders from capturing simplistic and non-informative trivial representations. This approach allows us to harness GNN encoder to learn high-level semantic representations while maintaining nearly the same computational demand during back-propagation.
IV-B Block Masking Process for Masked Autoencoding

### v2

@@ Abstract
Graph Neural Networks (GNNs) have shown promise in learning dynamic functional connectivity for distinguishing phenotypes from human brain networks. However, collecting large-scale labeled clinical data is often resource-intensive, limiting their practical use. This motivates the use of self-supervised learning to leverage abundant unlabeled data. While generative self-supervised approaches—particularly masked autoencoders—have shown promising results in representation learning in various domains, their application to dynamic graphs for dynamic functional connectivity remains underexplored and faces challenges in capturing high-level semantic patterns evolving over time. Here, we propose the Spatio-Temporal Joint Embedding Masked Autoencoder (ST-JEMA), which introduces a dual reconstruction objective that integrates spatial and temporal signals in dynamic brain graphs, enabling the encoder to learn temporally-aware semantic representations from unlabeled fMRI data. This design addresses key challenges in fMRI representation learning by capturing the spatio-temporal structure of dynamic functional connectivity. Unlike prior approaches that focus solely on static or spatial components, our method jointly reconstructs temporal dynamics and structural variations across time. Utilizing the large-scale UK Biobank dataset for self-supervised learning, ST-JEMA demonstrates robust representation learning performance across eight benchmark resting-state fMRI datasets, outperforming prior self-supervised methods in phenotype prediction and psychiatric diagnosis tasks—even with limited labeled data. Moreover, ST-JEMA proves resilient under temporal missing data scenarios, emphasizing the value of temporal reconstruction objectives. These findings demonstrate the effectiveness of our method for learning robust representations from fMRI data, particularly in label-scarce scenarios.
Keywords Functional connectivity  $\cdot$ Dynamic brain networks  $\cdot$ Spatiotemporal graph  $\cdot$ Self-supervised learning  $\cdot$ Representation learning
1 Introduction
Recent advancements in neuroimaging research have shown a growing interest in leveraging functional Magnetic Resonance Imaging (fMRI), which captures the Blood-Oxygen-Level-Dependent (BOLD) signal to understand the intricate functionalities of the human brain [1, 2]. Consequently, leveraging fMRI data through Functional Connectivity (FC) has gained traction in addressing various human brain-related problems using neural network models [3]. In this context, FC refers to the temporal correlation degree between Regions of Interest (ROI) of the brain, enabling the mathematical interpretation of fMRI data as graph structures, where nodes represent ROIs and edges represent the connectivity or correlation between these regions. This enables the utilization of graph-based learning methods such as Graph Neural Network (GNN) to analyze and interpret brain dynamics [4, 5].
@@ 1 Introduction
In response, utilizing unlabeled fMRI data through Self-Supervised Learning (SSL) presents a promising solution to extract valuable features for downstream tasks [6, 7, 8]. SSL, focused on pre-training backbone models without labels, has driven performance enhancements in diverse domains, including Natural Language Processing (NLP) [9], Image [10, 11], and Video Processing [12]. Among various SSL paradigms, generative approaches inspired by Masked Language Modeling (MLM) have shown superior performance compared to the contrastive learning methods in downstream tasks [11, 12] across a range of different tasks.
@@
This trend has also emerged in graph domains, as demonstrated by GraphMAE [8], which applies a similar philosophy to static graphs and outperforms traditional contrastive SSL approaches. In the context of dynamic graphs, Choi et al. [13] have developed a generative SSL framework that enables foundational models to effectively capture temporal dynamics. When applied to fMRI datasets, this framework demonstrates enhanced performance across multiple downstream tasks, surpassing models based on static graph methodologies and contrastive learning approaches.
@@
While recent advancements in generative SSL have shown improved performance in learning graph data structures, existing methods tend to focus on reconstructing lower-level features [8, 13], potentially limiting their ability to grasp higher-level semantic representations. Focusing on low-level features hinders the ability of the model to capture generalizable representations for downstream tasks.
@@
To tackle this limitation of current generative SSL methods — reconstructing low-level representation in temporal graph data — we propose a novel framework called Spatio-Temporal Joint Embedding Masked Autoencoder (ST-JEMA) for dynamic graphs such as dynamic functional connectivity of fMRI data. In ST-JEMA, we propose a new joint embedding reconstruction loss based on the latest SSL frameworks known as Joint-Embedding Predictive Architecture (JEPA) [14]. The JEPA-inspired loss encourages the encoder to reconstruct latent representations rather than raw features, incorporating both spatial and temporal dimensions. This encourages the GNN models to learn more advanced representations of dynamic graphs.
@@
We employed the extensive UK Biobank fMRI dataset, which consists of 40 913 records, for comparing various SSL methods without demographic characteristics (e.g. Gender), and validated our methodology by fine-tuning across eight benchmarks including non-clinical and clinical fMRI datasets. Experiments showed that ST-JEMA effectively improves node representation learning compared to baseline SSL methods. This improvement was evident in its superior performance on downstream fMRI tasks.
@@
  • We propose ST-JEMA, a novel framework that shifts generative SSL from low-level feature reconstruction toward high-level representation learning in dynamic graphs.
@@
Figure 1: Overview of the Spatio-Temporal Joint Embedding Masked Autoencoder (ST-JEMA) framework. (a) Pre-training pipeline, where the model jointly reconstructs node representations and edge structures by masking blocks in both spatial and temporal dimensions. The context and target GNN encoders are updated via EMA. Spatial and temporal processes are indicated by green and orange dashed lines, respectively. (b) Fine-tuning pipeline for downstream tasks (classification or regression), where the pre-trained encoder outputs are pooled across spatial and temporal dimensions to generate task-specific representations.
2 Related Works
2.1 Learning Functional Connectivity of the Brain with Graph Neural Networks
Recent advances in GNNs have facilitated the modeling of brain networks using FC derived from fMRI data. Pioneering GNN-based approaches tackle the high-dimensional nature of fMRI by modeling brain dynamics [15], employing ROI-aware architectures [16], or introducing generative frameworks [17].
@@ 2.1 Learning Functional Connectivity of the Brain with Graph Neural Networks
2.2 Self-supervised Learning on Graph Data
SSL effectively extracts meaningful representations from large-scale unlabeled graph data. In the realm of graph contrastive SSL, Deep Graph Infomax [20] and GraphCL [6] have been pioneering efforts that leverage augmentations and multiple graph views to enrich representation learning. However, challenges in augmentation efficacy across domains have led to innovative works, which simplify the process by introducing perturbations to embeddings [21] or model parameters [7]. Generative approaches in graph self-supervised learning focus on reconstructing the original graph structure or node features from encoded representations to capture inherent structural relationships. VGAE [22] employs variational autoencoders for adjacency matrix reconstruction enhancing link prediction performance, while GraphMAE [8] employs a masked autoencoder approach, which shows enhanced performance in node and graph classification compared to previous contrastive methods.
@@ 2.2 Self-supervised Learning on Graph Data
2.3 Joint Embedding Predictive Architecture
@@
In the evolving landscape of SSL in computer vision, Assran et al. [14] introduces the JEPA, a new framework that integrates the concepts of both Joint Embedding Architecture (JEA) in contrastive learning and generative learning process of the Masked Autoencoder [MAE; 11]. This approach diverges from the MAE, which mainly focuses on pixel-level image reconstruction. Unlike MAE, which reconstructs masked patches, JEPA aims to avoid the potential pitfall of generating representations overly fixated on low-level details. Instead, Assran et al. [14] utilizes a predictive model trained to infer the representations of masked target patches based on unmasked context patches, enabling the encoder to formulate representations that encapsulate higher-level semantic information beyond mere pixel details.
Refer to Appendix G for more detailed discussions on the related works.
3 Preliminaries
3.1 Problem Definition
We aim to develop a SSL framework for capturing temporal dynamics in dynamic graphs constructed from large-scale, unlabeled fMRI data. We pre-train a GNN encoder $f(\cdot;{\boldsymbol{\theta}})$ and decoder $g(\cdot;{\boldsymbol{\phi}})$ using our SSL strategies, and discard the decoder during downstream fine-tuning with prediction heads $h(\cdot;{\boldsymbol{\omega}})$. The objective is to obtain an encoder $f(\cdot;{\boldsymbol{\theta}}^{*})$ that captures transferable representations captured the key patterns and dynamics within the fMRI data under limited supervision.
3.2 Dynamic Graph Construction from fMRI data
We represent fMRI-based brain activity as a dynamic graph sequence ${\mathcal{G}}(t)=({\mathcal{V}}(t),{\mathcal{E}}(t))$ over discrete time $t\in{\mathcal{T}}$. Each node $i\in{\mathcal{V}}(t)$ corresponds to a brain region (ROI) defined by the Schaefer atlas [23], with $N$ representing the number of ROIs. Each node is associated with a feature vector $\boldsymbol{x}_{i}(t)\in\mathbb{R}^{d_{v}}$. The node feature is defined as:
where $\boldsymbol{e}_{i}$ is a spatial embedding and ${\boldsymbol{\eta}}(t)$ is a temporal encoding from a GRU [24] module.
Edges between nodes are defined using functional connectivity computed via the Pearson correlation of BOLD signals within a sliding temporal window:
@@ 3.2 Dynamic Graph Construction from fMRI data
3.3 Context and Target for Reconstruction
Our method expands generative SSL by adapting the core idea of JEPA [14]—reconstructing latent representations instead of input features—into the domain of dynamic graphs derived from fMRI data. While JEPA operates on static image patches, we introduce a multi-block masking strategy suited to spatio-temporal brain graphs. Each graph ${\mathcal{G}}(t)$ is partially masked with $K$ different binary masks, enabling learning from diverse context-target configurations.
Specifically, we define context and target node sets at time $t$ as:
@@ 3.3 Context and Target for Reconstruction
where $\boldsymbol{M}^{(k)}_{\boldsymbol{X}}(t)\in\{0,1\}^{N\times d_{v}}$ is the $k$-th binary mask.
In parallel, we apply the same masking strategy to the adjacency matrices $\boldsymbol{A}(t)$, where masked edges are excluded from message passing and subsequently reconstructed using the contextual node embeddings. This allows the model to jointly learn node-level and connectivity-level representations in the latent space.
4 Methods
4.1 Joint Embedding Architecture
@@
Employing separate encoders for context and target components is crucial to prevent representation collapse [25] during training [14]. Although these encoders share identical structural designs, they follow distinct parameter update rules. Specifically, the context encoder $f_{\texttt{cxt}}$ is updated via Stochastic Gradient Descent (SGD), driven by subsequent reconstruction loss objectives, while the target encoder $f_{\texttt{tar}}$ remains frozen.
Subsequently, The parameters of target encoder, $\bar{{\boldsymbol{\theta}}}$, are then updated using Exponential Moving Average (EMA) based on the evolving parameters of $f_{\texttt{cxt}}$ as follows:
@@
where hyperparameter $\beta\in[0,1]$ represents the decay factor. Here, $\bar{{\boldsymbol{\theta}}}_{\text{old}}$ and $\bar{{\boldsymbol{\theta}}}_{\text{new}}$ refer to the parameters of $f_{\texttt{tar}}$ before and after applying EMA process, respectively, and ${\boldsymbol{\theta}}_{\text{ctx}}$ indicates the current parameters of $f_{\texttt{cxt}}$. Adopting distinct update rules for each encoder helps avoid simplistic and non-informative trivial representations. This approach allows us to harness GNN encoder to learn high-level semantic representations while maintaining nearly the same computational demand during back-propagation.
4.2 Block Masking Process for Masked Autoencoding